OpenStudy (pottersheep):

I really need help with this Trig identity please!(Grade 12!) How do I prove that sinx-siny = 2cos([x+y]/2)sin([x+y]/2)

OpenStudy (pottersheep):

I've been at it for so long! I cant get it :(

OpenStudy (anonymous):

I can help.

OpenStudy (pottersheep):

Thank you so much

OpenStudy (pottersheep):

We have only learnt compound angles, by the way

OpenStudy (anonymous):

So there is an identity that's called co function identies, do you know them?

OpenStudy (pottersheep):

I think so

OpenStudy (anonymous):

Or sum to product identities. Do any of these sound familiar.

OpenStudy (pottersheep):

Is it the cos(90-x) kind of thing?

OpenStudy (anonymous):

um, not really. Here, let me post a link.

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (pottersheep):

Thanks Ill read it!

OpenStudy (pottersheep):

nope....never learnt those :s

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Give me like 5-10 minutes. This one is a little tricky. I'll try to keep you updated with the steps.

OpenStudy (pottersheep):

Alright I really appreciate it

OpenStudy (pottersheep):

I found a note that our teacher wrote on the board! I dont get it though... he wrote

OpenStudy (pottersheep):

sinx + siny = sin(a+b) = sinacosB + cosAsinB sin(a-b) = sinAcosB - cosAsinB sin(a+B) + sin (a - b) = 2sinAcosB

OpenStudy (pottersheep):

ignore the caps they are the same letter

OpenStudy (pottersheep):

Then he wrote : let : a+b = x a - b = y 2a = x+y x = (x+y)/2

OpenStudy (anonymous):

still working....

OpenStudy (anonymous):

Okay, Do you know the double angle formula?

OpenStudy (pottersheep):

yep!

OpenStudy (anonymous):

good.

OpenStudy (anonymous):

okay, so cos([x+y]/2)sin([x+y]/2). This equal sin(x+y)

OpenStudy (anonymous):

Are you following what i mean here?

OpenStudy (pottersheep):

hmm

OpenStudy (anonymous):

how about this. 2cosAsinA=sin2a. A=(x+y)/2

OpenStudy (pottersheep):

alright

OpenStudy (anonymous):

cosine of x is sine of (90-x) right?

OpenStudy (pottersheep):

yep

OpenStudy (anonymous):

okay. so =sin(x+y) =sinxcosy+sinycosx =sinx sin(90-y)+siny sin(90-x)

OpenStudy (pottersheep):

yep

OpenStudy (anonymous):

I have a quick question, is it sin(x+y)/2 or sin(x-y)/2?

OpenStudy (pottersheep):

one sec for the original question?

OpenStudy (anonymous):

yes.

OpenStudy (pottersheep):

oh no! you are right it sin(x-y)!!!!!!!

OpenStudy (pottersheep):

sorry ><

OpenStudy (anonymous):

Ok, that makes more sense now.

OpenStudy (anonymous):

Give me another minute to type this up.

OpenStudy (pottersheep):

Of course

OpenStudy (anonymous):

ok, i've got it.

OpenStudy (pottersheep):

:D

OpenStudy (anonymous):

ok let me know if you can see it.

OpenStudy (pottersheep):

I can see it! Im following through fine so far. Thank you so much

OpenStudy (anonymous):

Yep.